OpenStudy (anonymous):

A+B+C=180, then Σ tan(A/2)tan(B/2)= ?

OpenStudy (anonymous):

i'll post what all I know :

OpenStudy (anonymous):

A+B=180-C And.. We need to find - tan a/2 tan b/2 + tan b/2 tan c/2+ tan c/2 tan a/2

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (poopsiedoodle):

just a second let me do something for spanish class really quick and I'll take a look at it

OpenStudy (anonymous):

Thank you :) take your time :)

OpenStudy (poopsiedoodle):

Hmm... So, it A+2B = 180, and A+B = 180 - C, then B must equal C.

OpenStudy (poopsiedoodle):

And if that's the case, then A+2C = 180, and A+C = 180 - C.

OpenStudy (anonymous):

Oh i'm extremely sorry, it ain't a+b+b, it's a+b+c. .. M mistake

OpenStudy (poopsiedoodle):

oh lol. Just a second then

OpenStudy (anonymous):

lol

OpenStudy (poopsiedoodle):

A + B = 180 - C. So, I guess we're going to try to make some substitution happen here. A = 180 - C - B.

OpenStudy (poopsiedoodle):

So, 180 - C - B + B = 180 - C.

OpenStudy (poopsiedoodle):

Aaaaaaaaaaaaand that gets us nowhere, because after simplifying, we get 180 - C = 180 - C. Great.

OpenStudy (anonymous):

Haha..

OpenStudy (anonymous):

What if we convert tan in terms of sin's and cos' ?

OpenStudy (poopsiedoodle):

In that case, it looks like I don't quite know what I'm doing here. Soooooo, @mathmale. And I dunno. Let's try that, I guess. This is a triangle we're working with here, right?

OpenStudy (anonymous):

i tried doing it that way, but it doesn't really took me anywhere.. :(

OpenStudy (anonymous):

Yep. Right.

OpenStudy (poopsiedoodle):

Any side lengths given?

OpenStudy (anonymous):

didn't *

OpenStudy (anonymous):

Nopes. Just what i wrote in the question.

OpenStudy (poopsiedoodle):

Well dag, yo.

OpenStudy (poopsiedoodle):

Well, I dunno, it looks like 2 might be a side length, because tan(A/2) and such. When you do tangents, you divide the opposite side by the adjacent side. So, one of the sides is probs 2.

OpenStudy (anonymous):

No no .. i'm afraid that isn't the way we do it :/ Thanks for trying though! :D Thank you!

OpenStudy (poopsiedoodle):

Lel. Yeah, I'm not too good with geometry. But I tried :p And looks like we have hartnn here now :D

hartnn (hartnn):

if A+B+C = pi what does A/2 + B/2 equal ?

hartnn (hartnn):

i am intending to use \(\tan (x+y)\) formula

OpenStudy (anonymous):

told ya already - pi/2- c/2

OpenStudy (anonymous):

I tried it that way too.. but then it just looked so complicated!

hartnn (hartnn):

you mean pi/2 - c/2 ? \(\tan (\pi/2 -C/2) = (1/\tan C/2)\) now expand \(\large \tan (A/2+B/2)\) using the formula, what u get ?

hartnn (hartnn):

did u get how its 1/tan (C/2) ?

OpenStudy (anonymous):

tan[a/2]+tan[b/2]/ 1-tan a/2 tan b/2 = cot c/2 = 1/tanc/2

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

multiply tan c/2 on both sides :)

hartnn (hartnn):

now its just some re-arrangement of terms

OpenStudy (anonymous):

@hartnn - Oh my goodness.. Man, thank you. Thank you so very much! Thanks a ton! :)))

hartnn (hartnn):

you're most welcome ^_^

OpenStudy (anonymous):

Seriously, THANKS A TON!